OpenStudy (anonymous):

What percent of 60 is 12? Show your work.

OpenStudy (sburchette):

Well, know that a percentage is obtained by multiplying a quantity (60) by a percentage (x). So we can set up an equation that states, \[60x=12\] Just divide by 60 and convert into percentage form.\[x=12/60=.2=20percent\]

OpenStudy (anonymous):

It's 20% of 60. 12 / 60 * 100 = 20

OpenStudy (anonymous):

You simply divide 12/60=1/5=.2 so 12 is 20% of 60 as you move the decimal over by two places. Also, to check your answer you should have .2X60=12 as this states "20 percent of 60 is 12". (Note "of" equals to "times")